the number of defective production in a production process follows a poisson distribution with a mean of 2.6 per month, for a given month what is the probability there will be fewer than two defective production?


1
Expert's answer
2022-02-15T13:52:43-0500

Let the random variable "X" represent the number of defective productions in a production process the "X\sim Poisson(\lambda=2.6)" given as,

"p(X=x)={e^{-2.6}2.6^x\over x!}, \space x=0,1,2,3,..
\\0,\space elsewhere"

We find the probability,

"p(X\lt2)=p(X=0)+p(X=1)"

Now,

"p(X=0)={e^{-2.6}2.6^0\over 0!}=0.07427358" and "p(X=1)={e^{-2.6}2.6^1\over 1!}=0.19311131". Therefore, "p(X\lt 2)=0.07427358+0.19311131=0.26738489"

Thus, the probability that for a given month there will be fewer than two defective productions is 0.26738489